# Statistical learning theory and Occam’s razor: The argument from empirical risk minimization

### v1

@@ 4.2. The epistemology of machine learning methods
Much of modern philosophy of science implicitly or explicitly views this general epistemological project as a dead end. Reichenbach, in the words of van Fraassen (2000, p. 254), thought that empiricist epistemology must reject “Rationalism’s stringent criterion of adequacy: that an epistemology must show how absolutely reliable knowledge is possible.” Van Fraassen (1989; 2000; 2004) himself offers an outlook of an empiricism in explicit rejection of “defensive epistemology” which “concentrates on justification, warrant for, and defence of one’s belief” (1989, p. 170). Peirce, in the words of Levi (1998, p. 177), “explicitly dismissed doxastic skepticism when he observed that merely writing down a question challenging some current assumption is not sufficient to create the sort of doubt that should occassion an inquiry.” Levi (1980; 2004) himself further develops Peirce’s doubt-belief model of inquiry in explicit rejection of “pedigree epistemology,” under which “one is obliged to justify current beliefs” (2004, p. 11). The broad alternative project that arises is a pragmatist one where we take seriously that one will always already start with a body of beliefs that one at that time does not actively or genuinely doubt. The interesting question is not whether one actually has an ultimate justification for these beliefs. The interesting question is how to proceed from these beliefs: how to improve these beliefs. This leads to a forwards-looking epistemological project that investigates how to improve (refine, revise, update) beliefs in the light of new data. And in this project there is still an interesting question of justification: we can theorize about better and worse ways of doing so. [^34] This broad epistemological approach is certainly not particular to the authors I sampled here. The repudiation of quests for absolute justification is found in the writings of many other prominent philosophers, like Popper (“[t]he piles are driven down from above into the swamp,” Popper, 2002/1959, p. 94); the “model-relativity” of all our knowledge is already central to Kant., [^35] Sober (2015, pp. 86f) also sets apart, with reference to Neurath’s boat, a “foundationalist” picture from a “more defensible” alternative picture where we “now have numerous beliefs about the world. Our task is to take new observations into account so as to improve our system of beliefs. We don’t start from zero; we start from where we are” (ibid., p. 87). Notably, though, he presents these two pictures as two different versions of Bayesianism; there is no reference to these different epistemological views in his discussion of frequentist ideas. Other authors, including Levi and van Fraassen, also assume a Bayesian framework. I do not do that here: I am not after a formal account of the wider epistemic context of agents using machine learning methods to inform their beliefs. I am here interested in the epistemological lessons we may draw from a formal account of the machine learning methods themselves, namely, machine learning theory.
@@
The same sentiment can be found in the machine learning literature. Already thirty years ago, Russell (1991, p. 45), after discussing the infeasibility of “tabula rasa” learning, writes that “the picture that is currently fashionable in machine learning is that of an agent that already knows something and is trying to learn some more.” In similar vein, Domingos (2012, p. 81) himself, after attributing to Hume and Wolpert the insight that “data alone is not enough,” writes that “induction (what learners do) is a knowledge lever: it turns a small amount of input knowledge into a large amount of output knowledge.” Importantly, these authors, like also Shalev-Shwartz and Ben-David (2014, p. 94), do not just point out the impossibility of inductive inference without assumptions. They presuppose that there is always a starting point of initial knowledge, and put to one side the question of the actual basis for this supposed knowledge. They instead take machine learning to be about the forwards-looking question of how to best proceed from initial knowledge: how to learn more, or turn input knowledge into more output knowledge.
@@
The model-relative guarantees derived within the theory of machine learning serve exactly such a perspective. Model-relative theoretical guarantees concern algorithms that presuppose the instantiation, in each application, of an inductive model, that codifies specific prior knowledge. And these guarantees show that such algorithms are good relative to the instantiated inductive model, relative to the prior knowledge. This fits a picture where any real-world learning problem arises in a context where we already take many things for granted, and are willing to accept as prior knowledge. Given our starting point (our particular learning problem and goal, and the way we codify prior knowledge in a formal inductive model), the relevant theoretical model-relative guarantee advises us on how to proceed (what model-dependent algorithm is good to use). Learning-theoretic results thus provide a normative component to a forwards-looking epistemological perspective on machine learning methods. [^36] Following up on footnote 30, prior knowledge is not only instantiated in the inductive model qua hypothesis class, but also already in the other formal components of the learning problem. But clearly learning-theoretic guarantees must always be relative to these choices as well.
@@ 4.2.2. The theory and the practice
In evoking a wider epistemological perspective on machine learning, we have a responsibility to do justice, not only to the normative role of the theory, but also to the actual practice of (modern) machine learning. Here arises a worry that the preceding picture is excessively neat. For one thing, it is not exactly standard machine learning practice to carefully specify an inductive model strictly on the basis of well-formulated domain-specific assumptions; the practice, say in deep learning, is to a significant extent one of trial-and-error of different general architectures and hyperparameters, that encode largely ill-understood inductive biases. For another, the kind of justification that practitioners offer is less based on the (formal) properties of the learning algorithm than on the empirical performance of the output classifier (the trained model), in the first instance on a separate test data set or in a cross-validation procedure. [^37] This turn away from explicit modeling and towards predictive performance is what is indeed often seen to separate machine learning from “traditional” statistical inference (Breiman, 2001). A full reply to this worry would have to engage more systematically with the relation between the theory and the practice of machine learning, which I do not attempt in this paper. Here I will just observe the following. Whatever the several back-and-forths of design and evaluation in an actual machine learning pipeline, the core remains the training of an algorithm to return a classifier that generalizes well. And if there is one theoretical lesson that will always stand, then it is that the algorithm must possess restrictive inductive biases, and that the algorithm will only do well if the inductive biases are appropriate. It is also nothing short of a practical necessity to constrain, amidst all the further guesswork and trial-and-error, the possible inductive models to at least some extent; and this will still always involve at least some amount of knowledge about what is likely to be appropriate in the current problem. Going further, we can make the minimal normative point that it is indeed better to try and implement inductive biases that are aligned with what we believe or are prepared to assume about the relevant domain. With all the qualifications and idealizations involved in linking the theory to the practice (which also ties back to the qualifications listed in section 4.1.1 above), if there is a normative role for the theory to play, then it is in grounding learning procedures that can capitalize on an inductive model encoding inductive assumptions, fitting in a broad forward-looking epistemological picture on machine learning methods.
@@
The core principle of formal learning theory, which also has its roots in machine learning theory, [^38] Specifically, the approach of algorithmic learning theory going back to Gold (1967); see Jain et al. (1999). is that inductive problems call for a context-dependent means-ends analysis of what epistemic notions of success (ends) are attainable with what assumptions and methods (means). Schulte (1999) therefore also speaks of “means-ends” epistemology.
This means-ends analysis is context-dependent in the sense that given a particular learning problem, which usually comes with restrictive “background assumptions,” the analysis does not question these assumptions (Kelly, 1996, p. 11; Schulte, 2017, sects. 1.3, 2.2). For a particular learning problem and notion of success, the analysis is concerned with showing that certain methods can or cannot “solve” the problem (can or cannot attain the notion of success), given the background assumptions. This again fits a forwards-looking epistemological perspective (cf. Kelly, 2016, pp. 713f), where the analysis provides a model-relative justification (with the inductive model constituted by the background assumptions) for methods that solve the problem.
@@
In the logical framework studied in formal learning theory, there arises a neat hierarchy where different notions of success are characterized by the topological structure of the problem and background assumptions (Kelly, 1996). In the statistical learning theory framework, the fundamental theorem characterizes the main notion of success in terms of the combinatorial structure of the background assumptions: for a method to attain the success guarantee of learnability, the hypothesis class needs to have finite VC dimension. Thus, adopting Kelly’s words, the fundamental theorem shows what sort of knowledge (form of hypothesis class) is necessary for reliable inductive inference (a learnability guarantee) to be possible. This provides a means-ends reason for modeling, if we can, our background assumptions in the form of a class of hypotheses with finite VC dimension, a simple class of hypotheses. From the combination of a forwards-looking perspective and a means-ends analysis thus arises a justification for preferring simplicity—albeit with important qualifications.
@@ 4.3.1. The argument
We face a certain problem of classification, that we are prepared to model as a problem in statistical learning. We enter this problem with further prior knowledge still; and we are interested in a method that is good relative to this prior knowledge. As a formalization of what it means for a method to be good relative to prior knowledge, we adopt the model-relative notion of learnability. Now the fundamental theorem tells us that for there to exist a method with this guarantee of learnability, we need to formulate a hypothesis class, as formalization of our prior knowledge, that is a VC class—that is simple. Only when the hypothesis class is simple, does there exist a method with the guarantee of learnability relative to this hypothesis class. This “transcendental deduction” gives us a means-ends justification for modeling, if we can, our prior knowledge in the shape of a simple class of hypotheses.
The previous reasoning was based on the black-and-white picture of learnability or no. The quantitative version of the fundamental theorem offers a more fine-grained version; but the essence of the argument is the same. Accepting the guarantee of learnability, we now recognize that stronger bounds give a stronger guarantee, and we take a method with a stronger guarantee to be better. Now the quantitative version of the fundamental theorem tells us that a VC class of lower VC dimension—a simpler hypothesis class—gives a stronger guarantee. This gives us a means-ends justification for modeling, to the extent we can, our prior knowledge in the shape of class that is maximally simple (of maximally low VC dimension).
@@ 4.3.2. Qualifications
But perhaps most importantly, the final step of the argument still comes with a crucial qualifier. We have a means-ends justification for modeling our prior knowledge in the form of a simpler class, to the extent we can. The theory can be seen to push into one direction, towards simplicity. However, there is also a pull towards the opposite direction. The actual context of the learning problem, and in particular our prior knowledge, generally pulls away from simplicity, in the direction of richer assumptions, of a more complex hypothesis class.
@@
Note that the prior knowledge we have in each instance is an absolutely crucial constraint in the argument. In the forwards-looking picture, we always enter a learning problem with informal prior knowledge (beliefs we have, assumptions we are willing to make), and it is this informal prior knowledge that we need to encode in some formal hypothesis class. There will always be some leeway here, simply because it will be a significant translation step from our informal knowledge to the formal object of a hypothesis class, but it is still a constraint. It will never be reasonable, for instance, to adopt a singleton hypothesis class (a maximally simple class of VC dimension 0), because that would mean there would not be a learning problem to begin with. Without this constraint by informal prior knowledge, the argument would lose the connection to the epistemic context of an actual learning problem, and collapse into the useless advice to always choose a class of VC dimension 0.
@@
The downside is that there will be learning problems where this constraint, this pull towards complexity, is too strong for the simplicity argument to still be meaningful. There will be situations, in particular, where we cannot plausibly capture our prior knowledge into a class of VC dimension sufficiently small to still yield useful bounds. If such situations are the rule rather than the exception, then this is a serious restriction in the scope of the argument.
@@
The good news is that there is hope of generalizing the argument, to work with a much weaker kind of prior knowledge. Namely, in rough terms, we can simultaneously evaluate a (ranked) sequence of multiple VC classes, and accomodate the theoretical push towards simplicity by striking a certain trade-off between empirical error and VC dimension. In fact, this trade-off can be automated into a learning procedure. This is Vapnik’s second “inductive principle,” or the method of structural risk minimization.
4.3.3. The argument via structural risk minimization
@@
Indeed, rather than on uniform convergence and ERM, discussions of Occam’s razor in statistical learning theory tend to focus on this method (Harman and Kulkarni, 2007, ch. 3; Kelly, 2008, p. 335; Shalev-Shwartz and Ben-David, 2014, sect. 7.3; Bargagli Stoffi et al., 2022). Moreover, this method is directly applicable to the problem of model selection, which takes center stage in the modern philosophical debate on Occam’s razor (Forster and Sober, 1994; Sober, 2015). Very briefly, the formal route to the method of structural risk minimization (SRM) is the following.
@@
We start with a generalized version of uniform convergence (definition 3). This generalized version applies to a weighted countable sequence of VC classes, and gives a uniform accuracy bound for each hypothesis that depends on the class that the hypothesis is in (Shalev-Shwartz and Ben-David, 2014, thrm. 7.4). Next, similar to how the ERM method is defined as minimizing a uniform convergence bound (section 2.4 above), the SRM method is defined by minimizing a generalized uniform convergence bound (ibid., p. 62). Finally, we can prove a certain weaker guarantee of nonuniform learnability for SRM (ibid., def. 7.1, thrm. 7.5).
I think that these formal ingredients form the basis for a generalized means-ends model-relative justification for simplicity, incorporating the argument from ERM developed here. However, the details of this generalized argument, and how they relate to existing discussions of SRM and simplicity in the literature, must wait for a follow-up paper. [^39] As mentioned in footnote 12, my argument and in particular the notion of justification is built on the epistemic end of learnability, disregarding the epistemic end of model assessment. Typically, though, the training or model fitting is the earlier step in a machine learning application, followed implicitly (by noting training error) or explicitly (by estimating performance with a test set) by model assessment, and possibly still an intermediate model selection (Hastie et al., 2009, ch. 7; Shalev-Shwartz and Ben-David, 2014, ch. 14). The argument in this paper can thus be seen to be restricted to the earlier step; but an extended argument based on SRM should also move beyond.

### v2

@@ 4.2. Epistemology and machine learning theory
Much of modern philosophy of science implicitly or explicitly views this general epistemological project as a dead end. Reichenbach, in the words of van Fraassen (2000, p. 254), thought that empiricist epistemology must reject “Rationalism’s stringent criterion of adequacy: that an epistemology must show how absolutely reliable knowledge is possible.” Van Fraassen (1989; 2000; 2004) himself offers an outlook of an empiricism in explicit rejection of “defensive epistemology” which “concentrates on justification, warrant for, and defence of one’s belief” (1989, p. 170). Peirce, in the words of Levi (1998, p. 177), “explicitly dismissed doxastic skepticism when he observed that merely writing down a question challenging some current assumption is not sufficient to create the sort of doubt that should occassion an inquiry.” Levi (1980; 2004) himself further develops Peirce’s doubt-belief model of inquiry in explicit rejection of “pedigree epistemology,” under which “one is obliged to justify current beliefs” (2004, p. 11). The broad alternative project that arises is a pragmatist one where we take seriously that one will always already start with a body of beliefs that one at that time does not actively or genuinely doubt. The interesting question is not whether one actually has an ultimate justification for these beliefs. The interesting question is how to proceed from these beliefs: how to improve these beliefs. This leads to an epistemological project that investigates how to improve (refine, revise, update) beliefs in the light of new data. And in this project there is still an interesting question of justification: we can theorize about better and worse ways of doing so. [^34] This broad epistemological approach is certainly not particular to the authors I sampled here. The repudiation of quests for absolute justification is found in the writings of many other prominent philosophers, like Popper (“[t]he piles are driven down from above into the swamp,” Popper, 2002/1959, p. 94); the “model-relativity” of all our knowledge is not just central to modern authors on induction like Howson (2000), but already to Carnap and indeed to Kant., [^35] Sober (2015, pp. 86f) also sets apart, with reference to Neurath’s boat, a “foundationalist” picture from a “more defensible” alternative picture where we “now have numerous beliefs about the world. Our task is to take new observations into account so as to improve our system of beliefs. We don’t start from zero; we start from where we are” (ibid., p. 87). Notably, though, he presents these two pictures as two different versions of Bayesianism; there is no reference to these different epistemological views in his discussion of frequentist ideas. Other authors, including Levi and van Fraassen, also assume a Bayesian framework. I do not do that here: I am not after a formal account of the wider epistemic context of agents using machine learning methods to inform their beliefs. I am here interested in the epistemological lessons we may draw from a formal account of the machine learning methods themselves, namely, machine learning theory.
@@
The same sentiment can be found in the machine learning literature. Already thirty years ago, Russell (1991, p. 45), after discussing the infeasibility of “tabula rasa” learning, writes that “the picture that is currently fashionable in machine learning is that of an agent that already knows something and is trying to learn some more.” In similar vein, Domingos (2012, p. 81) himself, after attributing to Hume and Wolpert the insight that “data alone is not enough,” writes that “induction (what learners do) is a knowledge lever: it turns a small amount of input knowledge into a large amount of output knowledge.” Importantly, these authors, like also Shalev-Shwartz and Ben-David (2014, p. 94), do not just point out the impossibility of inductive inference without assumptions. They presuppose that there is always a starting point of initial knowledge, and put to one side the question of the actual basis for this supposed knowledge. They instead take machine learning to be about how to best proceed from initial knowledge: how to learn more, or turn input knowledge into more output knowledge.
@@
The model-relative guarantees derived within the theory of machine learning serve exactly such a perspective. Model-relative guarantees concern algorithms that presuppose the instantiation, in each application, of an inductive model, that codifies specific prior knowledge. And these guarantees show that such algorithms are good relative to the instantiated inductive model, relative to the prior knowledge. This fits a picture where any real-world learning problem arises in a context where we already take many things for granted, and are willing to accept as prior knowledge. Given our starting point (our particular learning problem and goal, and the way we codify prior knowledge in a formal inductive model), the relevant theoretical model-relative guarantee advises us on how to proceed (what model-dependent algorithm to use). Learning-theoretic results thus provide a normative component to this general epistemological perspective on learning methods. [^36] Following up on footnote 30, prior knowledge is not only instantiated in the inductive model qua hypothesis class, but also already in the other formal components of the learning problem. But clearly learning-theoretic guarantees must always be relative to these choices as well.
@@ 4.2.2. The theory and the practice
In evoking a wider epistemological perspective on machine learning, we have a responsibility to do justice, not only to the normative role of the theory, but also to the actual practice of (modern) machine learning. Here arises a worry that the preceding picture is excessively neat. For one thing, it is not exactly standard machine learning practice to carefully specify an inductive model strictly on the basis of well-formulated domain-specific assumptions; the practice, say in deep learning, is to a significant extent one of trial-and-error of different general architectures and hyperparameters, that codify largely ill-understood inductive biases. For another, the kind of justification that practitioners offer is less based on the (formal) properties of the learning algorithm than on the empirical performance of the output classifier (the trained model), in the first instance on a separate test data set or in a cross-validation procedure. [^37] This turn away from explicit modeling and towards predictive performance is what is indeed often seen to separate machine learning from “traditional” statistical inference (Breiman, 2001).
A full reply to this worry would have to engage more systematically with the relation between the theory and the practice of machine learning, which I do not attempt in this paper. Here I will just observe the following. Whatever the several back-and-forths of design and evaluation in an actual machine learning pipeline, the core remains the training of an algorithm to return a classifier that generalizes well. And if there is one theoretical lesson that will always stand, then it is that the algorithm must possess restrictive inductive biases, and that the algorithm will only do well if the inductive biases are appropriate. It is also nothing short of a practical necessity to constrain, amidst all the further guesswork and trial-and-error, the possible inductive models to at least some extent; and this will still always involve at least some amount of knowledge about what is likely to be appropriate in the current problem. Going further, we can make the minimal normative point that it is indeed better to try and implement inductive biases that are aligned with what we believe or are prepared to assume about the relevant domain. With all the qualifications and idealizations involved in linking the theory to the practice (which also ties back to the qualifications listed in section 4.1.1 above), if there is a normative role for the theory to play, then it is in grounding learning procedures that can capitalize on an inductive model encoding inductive assumptions, fitting in a broad “model-relative” epistemological picture on machine learning methods.
@@
The core principle of formal learning theory, which also has its roots in machine learning theory, [^38] Specifically, the approach of algorithmic learning theory going back to Putnam (1965) and Gold (1967); see Jain et al. (1999). is that inductive problems call for a context-dependent means-ends analysis of what epistemic notions of success (ends) are attainable with what assumptions and methods (means). Schulte (1999) therefore also speaks of “means-ends” epistemology.
This means-ends analysis is context-dependent in the sense that given a particular learning problem, which usually comes with restrictive “background assumptions,” the analysis does not question these assumptions (Kelly, 1996, p. 11; Schulte, 2017, sects. 1.3, 2.2). For a particular learning problem and notion of success, the analysis is concerned with showing that certain methods can or cannot “solve” the problem (can or cannot attain the notion of success), given the background assumptions. This again fits a “model-relative” epistemological perspective (cf. Kelly, 2016, pp. 713f), where the analysis provides a model-relative justification (with the inductive model constituted by the background assumptions) for methods that solve the problem.
@@
In the logical framework studied in formal learning theory, there arises a neat hierarchy where different notions of success are characterized by the topological structure of the problem and background assumptions (Kelly, 1996). In the statistical learning theory framework, the fundamental theorem characterizes the main notion of success in terms of the combinatorial structure of the background assumptions: for a method to have the success guarantee of learnability, the hypothesis class must have finite VC dimension. Thus, adopting Kelly’s words, the fundamental theorem shows what sort of knowledge (form of hypothesis class) is necessary for reliable inductive inference (a learnability guarantee). This provides a means-ends reason for modelling, if we can, our background assumptions in the form of a class of hypotheses with finite VC dimension, a simple class of hypotheses. From the combination of a model-relative perspective and a means-ends analysis thus arises a justification for preferring simplicity—albeit with important qualifications.
@@ 4.3. A qualified justification for simplicity
We can now assemble from the building blocks of the previous sections a justificatory argument.
@@ 4.3.1. The argument
We face a certain problem of classification, which we are prepared to model as a problem in statistical learning. We enter this problem with further prior knowledge still; and we are interested in a method that is good relative to this prior knowledge. As a formalization of what it means for a method to be good relative to prior knowledge, we adopt the model-relative notion of learnability. Now the fundamental theorem tells us that for there to exist a method with this guarantee of learnability, we need to formulate a hypothesis class, as formalization of our prior knowledge, that is a VC class—that is simple. Only when the hypothesis class is simple, does there exist a method with the guarantee of learnability relative to this hypothesis class. This “transcendental deduction” gives us a means-ends justification for modeling, if we can, our prior knowledge in the shape of a simple class of hypotheses.
The previous reasoning was based on the black-and-white picture of learnability or no. The quantitative version of the fundamental theorem offers a more fine-grained version; but the essence of the argument is the same. Accepting the guarantee of learnability, we recognize that stronger bounds give a stronger guarantee, and we take a method with a stronger guarantee to be better. Now the quantitative version of the fundamental theorem tells us that a VC class of lower VC dimension—a simpler hypothesis class—gives a stronger guarantee. This gives us a means-ends justification for modeling, to the extent we can, our prior knowledge in the shape of class that is maximally simple (of maximally low VC dimension).
@@ 4.3.2. Qualifications
But perhaps most importantly, the final step of the argument still comes with a crucial qualifier. We have a means-ends justification for modeling our prior knowledge in the form of a simpler class, to the extent we can. The theory can be seen to push into one direction, towards simplicity. However, the actual context of the learning problem, and in particular our prior knowledge, acts as a check on this push towards simplicity.
@@
The prior knowledge we have in each instance is a crucial constraint in the argument. In line with the epistemological “model-relative” perspective sketched above, we always enter a learning problem with informal prior knowledge (beliefs we have, assumptions we are willing to make), and it is this informal prior knowledge that we need to codify in some formal hypothesis class. There will always be some leeway here, simply because it will be a significant translation step from our informal knowledge to the formal object of a hypothesis class, but it is still a constraint. It will never be reasonable, for instance, to adopt a singleton hypothesis class (a maximally simple class of VC dimension 0), because that would mean there would not be a learning problem to begin with. Without this constraint by informal prior knowledge, the argument would lose the connection to the epistemic context of an actual learning problem, and collapse into the useless advice to always choose a class of VC dimension 0.
@@
The downside is that there will be learning problems where this constraint, this check from complexity, is too strong for the simplicity argument to still be meaningful. There will be situations, in particular, where our prior knowledge is too weak to plausibly translate into a class of VC dimension sufficiently small to still yield useful bounds. If such situations are the rule rather than the exception, then this seems to be a serious restriction in the scope of the argument.
@@
The good news is that we can embed the argument into a more general setting, presupposing a much weaker kind of prior knowledge. Namely, in rough terms, we can simultaneously evaluate a (ranked) sequence of multiple VC classes, where the theoretical push towards simplicity manifests itself again in a preference for lower-VC-dimension classes; but one which is now checked by the classes’ empirical error. In fact, this gives a trade-off which can be automated into a learning procedure. This is Vapnik’s second “inductive principle,” or the method of structural risk minimization.
4.3.3. Structural risk minimization
@@
Indeed, rather than on uniform convergence and ERM, discussions of Occam’s razor in statistical learning theory tend to focus on this method (Harman and Kulkarni, 2007, ch. 3; Kelly, 2008, p. 335; Shalev-Shwartz and Ben-David, 2014, sect. 7.3; Bargagli Stoffi et al., 2022). Moreover, this method is directly applicable to the problem of model selection, which takes center stage in the modern philosophical debate about Occam’s razor (Forster and Sober, 1994; Sober, 2015). Very briefly, the formal route to the method of structural risk minimization (SRM) is the following.
@@
We start with a generalization of uniform convergence (definition 3 above). This generalization applies to a weighted countable sequence of VC classes, and gives a uniform accuracy bound for each hypothesis which depends on the class that the hypothesis is in (Shalev-Shwartz and Ben-David, 2014, thrm. 7.4). Next, similar to how the ERM method is defined as minimizing a uniform convergence bound, which leads to a minimization of empirical risk (section 2.4 above), the SRM method is defined by minimizing a generalized uniform convergence bound, which leads to a minimization of a function of empirical risk and VC dimension (ibid., p. 62). This, finally, translates into a bound on SRM’s performance; and we can also prove a certain weaker guarantee of nonuniform learnability (ibid., def. 7.1, thrm. 7.5).
I think that SRM can be understood as implementing a parallel application, over multiple VC classes at the same time, of the core argument for simplicity that I gave above. The theoretical push towards simplicity is again a push towards hypotheses from classes of lower VC dimension. This push is again checked by the classes’ adequacy for the learning problem, or the adequacy of the inductive assumptions they codify. However, this is now not done by an informal evaluation of how well the formal inductive assumptions match our background knowledge. Instead, this adequacy is directly estimated by empirical error. This gives rise to a quantitative trade-off between these two elements, which SRM automates. By taking into account empirical errors, SRM can thus be seen to automate the evaluation of the adequacy of the inductive assumptions in individual VC classes; but of course the whole procedure is still relative to an initial choice of a weighted sequence of VC classes, which now constitutes the inductive model.
Clearly, it needs more work to spell out this view in the proper amount of detail, including how it fits with earlier discussions of Occam’s razor in SRM, and with the philosophical discussion of Occam’s razor in model selection methods. But that work must wait for another occasion.